. How many pints of a 9% cleaning solution must be mixed with 10 pints of a 13%
cleaning solution to give a 11% solution?
: . How many pints of a 9% cleaning solution must be mixed with 10 pints of a 13%
cleaning solution to give a 11% solution?

Answer by mangopeeler07(445)

x=pints of 9%
coefficient=percent

9x+13(10)=11(10+x)

Distribute
9x+130=110+11x

Subtract 9x from both sides
130=110+2x

Subtract 110 from both sides
20=2x

Divide both sides by 2
10=x

x=pints of 9%
------------------------------------------------------------------------------
Answer: 10 pints of 9%